# Conceptualization, Data Curation, Formal Analysis, Methodology, Software, Writing – original draft

### v1

@@ 3.3.1 Natural Language Prompt Embedding
To process natural language inputs, including prompts for general flight information, weather information, and NOTAMs, we utilize the tokenizer and embedding table provided by the selected LLM. We denote the tokenizer as a function $T(\cdot)$, which transforms each combined prompt $P_{i,t}$, where $P_{i,t}=P^{\text{fp}}_{i,t}\,\|\,P^{\text{wx}}_{t}\,\|\,P^{\text{no}}_{t},$ into a sequence of subword tokens. The resulting tokenized sequence is $T(P_{i,t})=[p_{1},p_{2},\dots,p_{L}],$ where $L$ denotes the length of the token sequence. Each token $p_{l}\in T(P_{i,t})$ is then mapped to a continuous embedding vector $\mathbf{z}^{p}_{l}\in\mathbb{R}^{d}$ using the pretrained embedding table $\text{Emb}\in\mathbb{R}^{V\times d}$, where $V$ is the vocabulary size and $d$ is the embedding dimension. The output of this step is the embedded prompt representation:
@@
where $\mathbf{Z}^{p}_{i,t}\in\mathbb{R}^{L\times d}$ is the token-level embedding sequence for prompt $P_{i,t}$. The embedding table and tokenizer are kept fixed during training to preserve the pretrained language understanding, ensuring that natural language inputs are represented consistently and remain aligned with the pretrained LLM.
@@ 3.3.2 Trajectory Representation Model
The three trajectory types, focusing $X^{f}_{i,t}$, active $X^{a}_{i,t}$, and prior $X^{p}_{i,t}$, are time-series data compressed into instance-level representations using a pretrained trajectory encoder, where each embedding represents one trajectory. Thus, it is crucial that each embedding preserves the semantic fidelity of its corresponding trajectories and aligns well with human understanding of air traffic patterns. ATSCC [56] is well-suited for air traffic trajectory data, showing strong performance in instance-level tasks like classification and clustering, with representations that align with aeronautical procedures without relying on predefined labels. Therefore, the trajectory representation encoder model $f_{atscc}(\cdot)$ follows the architecture and was pretrained in accordance with the reproduction details of ATSCC.
We reproduce the trajectory encoder by strictly following the reproduction details of the ATSCC framework [56]. The causal transformer encoder architecture features 12 layers with a model dimension of 768, feed-forward network dimensions of 3072 using the Gaussian Error Linear Unit (GELU) activation, and 12 attention heads. The model has an input projection from 9 to the model dimension and an output projection from the model dimension to 320. ATSCC follows the NoPos [82], which omits positional encoding. Additionally, L2 normalization is applied after both input and output projections. The model is pre-trained on ATFMTraj [83] data, using 5-second interval downsampling and combining both Incheon departures and arrivals. To enhance the expressiveness of trajectory data, additional geometric features, including direction vectors and polar coordinates, are computed. For each timestep $k$ in a time-series instance, the additional geometric features are calculated as:
$\{x_{i,k}^{u_{x}},x_{i,k}^{u_{y}},x_{i,k}^{u_{z}}\}$ $=\frac{x_{i,k+1}-x_{i,k}}{\left\|x_{i,k+1}-x_{i,k}\right\|},$ (15)
$\{x_{i,k}^{r},x_{i,k}^{\sin\theta},x_{i,k}^{\cos\theta}\}$ $=\left\{\sqrt{\left(x_{i,k}^{x}\right)^{2}+\left(x_{i,k}^{y}\right)^{2}},\,\sin\left(\arctan 2(x_{i,k}^{y},x_{i,k}^{x})\right),\,\cos\left(\arctan 2(x_{i,k}^{y},x_{i,k}^{x})\right)\right\}.$ (16)
$x_{k}$ $=\{x_{i,k}^{x},x_{i,k}^{y},x_{i,k}^{z},x_{i,k}^{u_{x}},x_{i,k}^{u_{y}},x_{i,k}^{u_{z}},x_{i,k}^{r},x_{i,k}^{\sin\theta},x_{i,k}^{\cos\theta}\}$ (17)
During pretraining, we apply a dropout rate of 0.35 and binomial input embedding masking with a probability of 0.2. The model is trained using the AdamW optimizer with a batch size of 16, a learning rate of $1\times 10^{-5}$, and weight decay of $1\times 10^{-5}$. A grid search is conducted by separating a small portion of the dataset for validation to determine the optimal segmentation threshold and contrastive loss temperature based on classification and clustering performance. After pretraining, the trajectory encoder weights are frozen and not updated further.
@@
The pre-trained encoder, denoted as $f_{\text{atscc}}(\cdot)$, first applies geometric extraction functions to $X^{(k)}_{i,t}\in\mathbb{R}^{N^{(k)}\times\mathcal{T}^{(k)}_{\max}\times 3}$ to expand each trajectory state into a 9-dimensional feature vector $X^{(k)}_{i,t}\in\mathbb{R}^{N^{(k)}\times\mathcal{T}^{(k)}_{\max}\times 9}$ with 5-second interval downsampling, followed by input projection and a forward pass through the causal transformer backbone. The final timestep output is then linearly projected to a 320-dimensional instance-level representation. The process of encoding the trajectory for type $k$ is illustrated in Figure 5 and formally defined as
@@
where $k\in\{f,a,p\}$ denotes the trajectory group and $Z^{(k)}_{i,t}=[z^{(k)}_{1},\dots,z^{(k)}_{N^{(k)}}]\in\mathbb{R}^{N^{(k)}\times 320}$ is the resulting representation. According to Figure 4, the trajectory representation encoder is placed at the beginning of the trajectory processing pipeline. Importantly, its weights are frozen throughout the training pipeline, meaning it does not participate in backpropagation. This design choice allows us to pre-encode the trajectories into their instance-level representations $Z^{(k)}_{i,t}$ prior to training. Doing so, we significantly improve memory efficiency, as we avoid training a highly memory-intensive nested transformer architecture, where long time-series sequences are first encoded into fixed-size vectors and then further embedded into a language-style sequence model. In contrast, pre-encoding the trajectories enables more efficient training without compromising the ability to incorporate rich temporal trajectory information.
@@ 3.3.3 Cross-modality Adaptation Network
To enable integration between language-based prompt embeddings and trajectory-based time-series representations, we employ a cross-modality adaptation network, denoted as $f_{\text{xa}}(\cdot)$, which adapts the trajectory inputs from the time-series modality into the language modality. This network is structured as a lightweight MLP, designed to project instance-level trajectory representations into the language model’s embedding space. Specifically, given the output from the pretrained trajectory encoder $Z^{(k)}_{i,t}\in\mathbb{R}^{N^{(k)}\times 320}$, the adaptation network transforms this vector into the LLM-friendly dimension $d$ as follows:
@@
where $\mathbf{Z}^{(k)}_{i,t}\in\mathbb{R}^{N^{(k)}\times d}$ is the LLM-compatible embedding of trajectory group $k\in\{f,a,p\}$ (focusing, active, or prior). The MLP architecture, which projects the 320-dimensional trajectory embedding to the LLM’s model dimension of $f_{\text{xa}}(\cdot)$, comprises two linear layers with GELU activation functions and a dropout rate of 0.35 applied after each activation. The cross-modality adaptation network is simple and efficient, introducing only a small computational cost while effectively bridging the gap between trajectory encoders and LLMs. Its ability to integrate with the LLM without architectural changes highlights the practicality of cross-modal integration with minimal engineering effort. The embeddings $\mathbf{Z}^{(k)}_{i,t}$ are inserted into designated positions within the tokenized static trajectory prompts, forming a unified input sequence of language and trajectory information processed by the LLM backbone. To combine the trajectories with textual embeddings, we insert trajectory embeddings between guiding prompts (Figure 12) to help the model distinguish different trajectory groups. The trajectory input is formally defined as:
$\mathbf{Z}^{T}_{i,t}=\text{Concat}(\text{Emb}[T(P_{st,1})],\ \mathbf{Z}^{f}_{i,t},\ \text{Emb}[T(P_{st,2})],\ \mathbf{Z}^{a}_{i,t},\ \text{Emb}[T(P_{st,3})],\ \mathbf{Z}^{p}_{i,t},\ \text{Emb}[T(P_{st,4})]),$ (20)
This concatenated sequence $\mathbf{Z}^{T}_{i,t}\in\mathbb{R}^{(L_{st,1}+1+L_{st,2}+N^{a}+L_{st,3}+N^{p}+L_{st,4})\times d}$ represents the complete trajectory-informed input for the multimodal model, where $L_{st,k}$ denotes the length of the $k$-th static prompt segment after tokenization and embedding, 1, $N^{a}$, and $N^{p}$ are added to the embedding sequence length, as we use the instance-level representation of each trajectory as a single token.
@@ 3.3.4 Large Language Model Backbone
To leverage pretrained linguistic understanding, we retain the LLM backbone while replacing its language modeling head with a regression head, enabling prediction of continuous post-terminal duration values. In this paper, we consider LLMs with sizes under 1.5 billion parameters, as the task is formulated as a single-value regression problem and does not require large-scale generative capacity. These small LLMs can run efficiently on a single consumer-grade GPU, demonstrating their suitability for deployment by a wider range of users. The candidate backbone models include LLaMA3.2-1B, LLaMA3.2-1B-Instruct, Qwen2.5-1.5B, Qwen2.5-1.5B-Instruct, DeepSeek-R1-Distill-Qwen-1.5B, TinyLlama-v1.1, TinyLlama-1.1B-Chat-v1.0, and Pythia-1B. All model parameters and the embedding table are kept frozen as provided by the original sources to preserve their pretrained linguistic capabilities and ensure memory efficiency.
The LLM backbone processes multimodal inputs consisting of textual prompt embeddings, $\mathbf{Z}^{p}_{i,t}$ obtained via the corresponding tokenizer and embedding table, and trajectory-informed embeddings $\mathbf{Z}^{T}_{i,t}$ generated by the representation and cross-modality networks. Therefore, the input sequence fed to the LLM backbone is formed by concatenating these two components:
@@ 3.4 Training Pipeline
For each scenario $s_{i,t}$, the textual components, general flight information, weather reports, and NOTAMs, are tokenized and embedded to form the prompt embedding $\mathbf{Z}^{p}_{i,t}$. Separately, the static prompts that describe the structure and semantics of the trajectory groups (focusing, active, and prior) are combined with their corresponding instance-level representations, which were pre-encoded by the frozen ATSCC model $f_{\text{atscc}}$ and projected via the trainable adapter $f_{\text{xa}}$ to form the trajectory-informed sequence $\mathbf{Z}^{T}_{i,t}$. Those embedding sequences are concatenated into a unified sequence $\mathbf{Z}_{i,t}$, which is then fed into the frozen LLM backbone $f_{\text{llm}}$. Since the LLM is a causal model, only the final hidden state $\mathbf{h}_{i,t}$ is extracted, representing a context-aware summary of the entire input sequence. This hidden state is passed to a trainable regression head $f_{h}$ to produce a scalar prediction $\hat{y}_{i}$, corresponding to the estimated time spent in the airspace. The ground-truth target $y_{i}$ is standardized prior to training. Optimization is performed using the AdamW optimizer with a learning rate of $1\times 10^{-5}$, weight decay of $1\times 10^{-5}$, and a mini-batch size of 4. The model is trained using the Smooth L1 Loss:
@@
which is robust to outliers, well-suited, and in some cases, prevents exploding gradients [84]. The parameters of the pretrained components, namely the embedding table Emb, LLM backbone $f_{\text{llm}}$, and the trajectory encoder $f_{\text{atscc}}$, are frozen. Only the cross-modality adaptation network $f_{\text{xa}}$ and the regression head $f_{h}$ are updated during training, as illustrated in Figure 4, ensuring memory efficiency by avoiding gradient allocation to the LLM and trajectory models. All training and evaluation are conducted using Python 3.11.8 and PyTorch 2.0.1 with CUDA Toolkit 11.8, running on an NVIDIA GeForce RTX 4090 GPU.
Algorithm 1 Multimodal Training Pipeline for Flight Delay Regression
@@
4:  $Z^{(j)}_{i,t}=f_{\text{atscc}}(X^{(j)}_{i,t}),\quad j\in\{f,a,p\}$
@@
11:    $\mathbf{Z}^{(j)}_{i,t}=f_{\text{xa}}(Z^{(j)}_{i,t}),\quad j\in\{f,a,p\}$
12:    $\mathbf{Z}^{T}_{i,t}=\text{Concat}(\text{Emb}[T(P_{st,1})],\ \mathbf{Z}^{f}_{i,t},\ \text{Emb}[T(P_{st,2})],\ \mathbf{Z}^{a}_{i,t},\ \text{Emb}[T(P_{st,3})],\ \mathbf{Z}^{p}_{i,t},\ \text{Emb}[T(P_{st,4})])$
@@
18:   Backpropagate and update trainable parameters in $f_{\text{xa}}(\cdot)$ and $f_{h}(\cdot)$

### v2

@@ 3.3.1 Natural Language Prompt Embedding
To process natural language inputs, including prompts for general flight information, weather information, and NOTAMs, we use the tokenizer and embedding table provided by the selected LLM. We denote the tokenizer as a function $T(\cdot)$, which transforms each combined prompt $P_{i,t}$, where $P_{i,t}=P^{fp}_{i,t}\,\|\,P^{wx}_{t}\,\|\,P^{no}_{t},$ into a sequence of subword tokens. The resulting tokenized sequence is $T(P_{i,t})=[p_{1},p_{2},\dots,p_{L}],$ where $L$ denotes the length of the token sequence. Each token $p_{l}\in T(P_{i,t})$ is then mapped to a continuous embedding vector $\mathbf{z}^{p}_{l}\in\mathbb{R}^{d}$ using the pretrained embedding table $\text{Emb}\in\mathbb{R}^{V\times d}$, where $V$ is the vocabulary size and $d$ is the embedding dimension. The output of this step is the embedded prompt representation:
@@
where $\mathbf{Z}^{p}_{i,t}\in\mathbb{R}^{L\times d}$ is the sequence of token embeddings for prompt $P_{i,t}$. The embedding table and tokenizer are kept fixed during training to preserve the pretrained language understanding, ensuring that textual inputs are represented consistently and remain aligned with the pretrained LLM.
@@ 3.3.2 Trajectory Representation Model
The three trajectory types, focusing $X^{f}_{i,t}$, active $X^{a}_{i,t}$, and prior $X^{p}_{i,t}$, as time-series data, are compressed into instance-level embeddings using a pretrained trajectory encoder, where each embedding represents one trajectory. Thus, it is crucial for each embedding to maintain the semantic integrity of its trajectory and align with human understanding of air traffic patterns. ATSCC [56] is well-suited for air traffic trajectory data, showing strong performance in classification and clustering tasks on the representations that align with aeronautical procedures without relying on predefined labels.
We reproduce the trajectory encoder $f_{\text{atscc}}(\cdot)$ by strictly following the reproduction details of the ATSCC framework. The causal transformer encoder architecture features 12 layers with a model dimension of 768, feed-forward network dimensions of 3072 using the Gaussian Error Linear Unit (GELU) activation, and 12 attention heads. The model has an input projection from 9 to the model dimension and an output projection from the model dimension to 320. ATSCC applies L2 normalization after both input and output projections and omits positional encoding [82]. The 9-dimensional input channel is designed to support a more expressive form of trajectory data by incorporating additional geometric features, including direction vectors and polar coordinates. For each timestep $\kappa$ in a time-series instance, the additional geometric features are calculated as:
$\{x_{i,\kappa}^{u_{x}},x_{i,\kappa}^{u_{y}},x_{i,\kappa}^{u_{z}}\}$ $=\frac{x_{i,\kappa+1}-x_{i,\kappa}}{\left\|x_{i,\kappa+1}-x_{i,\kappa}\right\|},$ (15)
$x_{i,\kappa}^{r}$ $=\sqrt{\left(x_{i,\kappa}^{x}\right)^{2}+\left(x_{i,\kappa}^{y}\right)^{2}}$ (16)
$x_{i,\kappa}^{\sin\theta}$ $=\sin\left(\arctan 2(x_{i,\kappa}^{y},x_{i,\kappa}^{x})\right)$ (17)
$x_{i,\kappa}^{\cos\theta}$ $=\cos\left(\arctan 2(x_{i,\kappa}^{y},x_{i,\kappa}^{x})\right)$ (18)
Therefore, at each time step $\kappa$, the trajectory feature is collected into the feature set
$x_{\kappa}=\{x_{i,\kappa}^{x},x_{i,\kappa}^{y},x_{i,\kappa}^{z},x_{i,\kappa}^{u_{x}},x_{i,\kappa}^{u_{y}},x_{i,\kappa}^{u_{z}},x_{i,\kappa}^{r},x_{i,\kappa}^{\sin\theta},x_{i,\kappa}^{\cos\theta}\}$ (19)
The model is pre-trained on ATFMTraj [83] data, using 5-second interval downsampling and combining both Incheon departures and arrivals. During pretraining, we apply a dropout rate of 0.35 and binomial input embedding masking with a probability of 0.2. The model is trained using the AdamW optimizer with a batch size of 16, a learning rate of $1\times 10^{-5}$, and weight decay of $1\times 10^{-5}$. A grid search is conducted by separating a small portion of the dataset for validation to determine the optimal segmentation threshold and contrastive loss temperature based on classification and clustering performance. After pretraining, the trajectory encoder weights are frozen and not updated further.
@@
The encoding process of the pre-trained encoder, denoted as $f_{\text{atscc}}(\cdot)$, first applies geometric extraction functions to $X^{(k)}_{i,t}\in\mathbb{R}^{N^{(k)}\times\mathcal{T}^{(k)}_{\max}\times 3}$ to expand each trajectory state into a 9-dimensional feature vector $X^{(k)}_{i,t}\in\mathbb{R}^{N^{(k)}\times\mathcal{T}^{(k)}_{\max}\times 9}$ with 5-second interval downsampling, followed by input projection and a forward pass through the causal transformer backbone. Only the final timestep output of each trajectory is then linearly projected to a 320-dimensional instance-level representation. The process of encoding the trajectory for type $k$ is illustrated in Figure 5 and formally defined as
@@
where $k\in\{f,a,p\}$ and $Z^{(k)}_{i,t}=[z^{(k)}_{1},\dots,z^{(k)}_{N^{(k)}}]\in\mathbb{R}^{N^{(k)}\times 320}$. According to Figure 4, the trajectory encoder is positioned at the start of the pipeline, with its weights frozen during training. This allows trajectories to be pre-encoded into instance-level representations $Z^{(k)}_{i,t}$ prior to training. This design significantly improves memory efficiency by avoiding a highly memory-intensive nested transformer training setup, where long time-series sequences are first encoded into fixed-size vectors and then reprocessed by another sequence model. In contrast, pre-encoding with the ATSCC encoder enables efficient training without compromising rich trajectory information.
@@ 3.3.3 Cross-modality Adaptation Network
To enable integration between textual prompt embeddings and trajectory representations, we employ a cross-modality adaptation network, denoted as $f_{\text{xa}}(\cdot)$. This network is structured as a lightweight MLP that projects instance-level trajectory representations into the language model’s embedding space, adapting time series into the language modality. Given the output from the pretrained trajectory encoder $Z^{(k)}_{i,t}\in\mathbb{R}^{N^{(k)}\times 320}$, the adaptation network transforms this vector into an LLM-compatible embedding with dimension $d$ as follows:
@@
where $\mathbf{Z}^{(k)}_{i,t}\in\mathbb{R}^{N^{(k)}\times d}$ is the embedding of trajectory group $k\in\{f,a,p\}$ (focusing, active, or prior). The MLP architecture consists of two linear layers with GELU activations and a dropout rate of 0.35 applied after each activation. The first layer maps the 320-dimensional embedding to the dimension of $d$, while the second layer preserves this dimension. The network is simple yet effective, introducing only a small computational cost while effectively bridging the gap between trajectory encoders and LLMs, highlighting the practicality of cross-modal adaptation with minimal engineering effort. The embeddings $\mathbf{Z}^{(k)}_{i,t}$ are inserted into designated positions within the tokenized static trajectory prompts, forming a unified input sequence of language and trajectory information processed by the LLM backbone. To combine trajectories with textual embeddings, we insert trajectory embeddings between the guiding prompts (Figure 12) to help the model distinguish between trajectory groups. The trajectory-informed sequence is denoted as:
$\mathbf{Z}^{T}_{i,t}=\text{Concat}($ $\text{Emb}[T(P_{st,1})],$
$\ \mathbf{Z}^{f}_{i,t},\ \text{Emb}[T(P_{st,2})],$
$\ \mathbf{Z}^{a}_{i,t},\ \text{Emb}[T(P_{st,3})],$
$\ \mathbf{Z}^{p}_{i,t},\ \text{Emb}[T(P_{st,4})]),$ (22)
The sequence $\mathbf{Z}^{T}_{i,t}\in\mathbb{R}^{(L_{st,1}+1+L_{st,2}+N^{a}+L_{st,3}+N^{p}+L_{st,4})\times d}$ represents the complete trajectory-informed seqeunce for the multimodal model, where $L_{st,k}$ denotes the length of the $k$-th static guiding prompt segment after tokenization and the number of embedding, 1, $N^{a}$, and $N^{p}$ are added to the embedding sequence length, as we represent each trajectory as a single token.
@@ 3.3.4 Large Language Model Backbone
To leverage pretrained linguistic understanding, we retain the LLM backbone while replacing its language modeling head with a regression head, enabling prediction of continuous post-terminal duration values. In this paper, we consider LLMs with sizes under 1.5 billion parameters, as the task is formulated as a single-value regression problem and does not require large-scale generative capacity. These small LLMs can run efficiently on a single consumer-grade GPU, demonstrating their suitability for deployment by a wider range of users. The candidate backbone models include LLaMA3.2-1B, LLaMA3.2-1B-Instruct, Qwen2.5-1.5B, Qwen2.5-1.5B-Instruct, DeepSeek-R1-Distill-Qwen-1.5B, TinyLlama-v1.1, TinyLlama-1.1B-Chat-v1.0, and Pythia-1B. All model parameters and the embedding table are kept frozen as provided by the original sources to preserve their pretrained linguistic capabilities and ensure memory efficiency. The LLM backbone processes multimodal inputs consisting of textual prompt embeddings, $\mathbf{Z}^{p}_{i,t}$ obtained via the corresponding tokenizer and embedding table, and trajectory-informed sequence $\mathbf{Z}^{T}_{i,t}$ generated by the representation and cross-modality networks. Therefore, the input sequence fed to the LLM backbone is formed by concatenating these two components:
@@ 3.4 Training Pipeline
For each scenario $s_{i,t}$, the textual components, including general flight information, weather reports, and NOTAMs, are tokenized and embedded to form the prompt embedding $\mathbf{Z}^{p}_{i,t}$. Separately, the static prompts that guide the LLM of the trajectory groups (focusing, active, and prior) are combined with their corresponding instance-level representations, which were pre-encoded by the frozen ATSCC model $f_{\text{atscc}}(\cdot)$ and projected via the trainable adapter $f_{\text{xa}}(\cdot)$ to form the trajectory-informed sequence $\mathbf{Z}^{T}_{i,t}$. Those embedding sequences are concatenated into a unified sequence $\mathbf{Z}_{i,t}$, which is then fed into the frozen LLM backbone $f_{\text{llm}}(\cdot)$. Since the LLM is a causal model, only the final hidden state $\mathbf{h}_{i,t}$ is extracted, representing a summary of the entire input sequence. This hidden state is passed to a trainable regression head $f_{h}(\cdot)$ to produce a scalar prediction $\hat{y}_{i}$, corresponding to the estimated time spent in the airspace. The ground-truth target $y_{i}$ is standardized prior to training. Optimization is performed using the AdamW optimizer with a learning rate of $1\times 10^{-5}$, weight decay of $1\times 10^{-5}$, and a mini-batch size of 4. The model is trained using the Smooth L1 Loss:
@@
which is robust to outliers, well-suited, and in some cases, prevents exploding gradients [84]. The parameters of the pretrained components, namely the embedding table Emb, LLM backbone $f_{\text{llm}}(\cdot)$, and the trajectory encoder $f_{\text{atscc}}(\cdot)$, are frozen. Only the cross-modality adaptation network $f_{\text{xa}}(\cdot)$ and the regression head $f_{h}(\cdot)$ are updated during training, as illustrated in Figure 4, avoiding memory allocation for the gradient of the LLM and the trajectory encoder. All training and evaluation are conducted using Python 3.11.8 and PyTorch 2.0.1 with CUDA Toolkit 11.8, running on an NVIDIA GeForce RTX 4090 GPU.
Algorithm 1 Training Pipeline for Flight Delay Regression
@@
4:  $Z^{(k)}_{i,t}=f_{\text{atscc}}(X^{(k)}_{i,t}),\quad k\in\{f,a,p\}$
@@
11:    $\mathbf{Z}^{(k)}_{i,t}=f_{\text{xa}}(Z^{(k)}_{i,t}),\quad k\in\{f,a,p\}$
12:    $\begin{aligned} \mathbf{Z}^{T}_{i,t}=\text{Concat}(&\text{Emb}[T(P_{st,1})],\\ &\mathbf{Z}^{f}_{i,t},\ \text{Emb}[T(P_{st,2})],\\ &\mathbf{Z}^{a}_{i,t},\ \text{Emb}[T(P_{st,3})],\\ &\mathbf{Z}^{p}_{i,t},\ \text{Emb}[T(P_{st,4})])\end{aligned}$
@@
18:   Backpropagate and update parameters in $f_{\text{xa}}(\cdot)$ and $f_{h}(\cdot)$